OpenStudy (anonymous):

I need help completing a table. ie y = x + 5: this is my slope-intercept form on the other side of this table I need to write the standard form for this equation. this is an example please show work.

OpenStudy (anonymous):

Ok.... y=x+5 So standard form is Ax+By=C

OpenStudy (anonymous):

all you need to do is subract x from both sides

OpenStudy (anonymous):

-x+y=5

OpenStudy (anonymous):

questions?

OpenStudy (anonymous):

can you show me how convert standard form to slope-intercept form ie 2x -3y = -6

OpenStudy (anonymous):

yes one sec

OpenStudy (anonymous):

Ok it's very simple let me explain... you'll get it in no time

OpenStudy (anonymous):

2x-3y=-6 (subtract 2x from each side) -3y=-2x-6 (divide each side by -3) y=2/3x+2